OpenStudy (anonymous):

Solve the equation. 42 = 3c

OpenStudy (anonymous):

c=42/3

OpenStudy (anonymous):

to isolate c divide both sides by 3

OpenStudy (anonymous):

14

OpenStudy (anonymous):

Correct! Thank you for the pretty medals!

OpenStudy (theviper):

42 = 3c c=42/3=14